On January 1, 2021, Drenz Co. acquired 30,000 ordinarily shares out of the 100,000 outstanding ordinary shares of Josiah Inc. for ₱5,000,000. Josiah’s assets and liabilities approximate their fair values except for inventories with carrying amount of ₱800,000 and fair value of ₱900,000 and machinery with carrying amount of ₱2,500,000 and fair value of ₱2,200,000. The remaining useful life of the machinery is 5 years. Josiah’s net asset has a book value of ₱10,000,000.


On December 31,2021. Josiah reported net income of ₱2,000,000 and declared and paid dividends of ₱800,000. On December 31,2022. Josiah reported net income of ₱4,500,000 and declared and paid dividends of ₱1,600,000.


5. How much is the carrying amount of the investment as of December 31,2022?


1
Expert's answer
2021-06-14T13:26:46-0400

Balance increasing:

"2,000,000-800,000+4,500,000-1,600,000-\frac{2,500,000}{5}*2=3,100,000"

Balance book values on On December 31, 2022:

"10,000,000+3,100,000=13,100,000"

Drenz Co. has 30% of Josiah Inc.

the carrying amount of the investment as of December 31,2022:

"13,100,000*0.3=3,930,000"